OpenStudy (anonymous):

a capacitor of 4 micro farad are connected in series with parallel combination of two capacitors of 6 micro farad each.The equivalent capacitance is ?

OpenStudy (anonymous):

Capacitors combine oppositely from resistors in parallel; C = C1 + C2 in series

OpenStudy (anonymous):

1/C = 1/C1 + 1/C2 so., combine the capacitors that are in parallel, then combine this result [C'] with the one in series, using the rules above.

OpenStudy (anonymous):

thankx